OpenStudy (anonymous):

In degrees, when does tangent equal 0?

OpenStudy (mathstudent55):

\(\tan \theta = \dfrac{\sin \theta}{\cos \theta} \) \(\tan \theta\) equals zero whrere \(\sin \theta\) equals zero.

OpenStudy (anonymous):

sin = 0 on the x axis which is 0 and 180

OpenStudy (danjs):

When the x value on the unit circle is Zero,